OpenStudy (anonymous):

Medal/Fan Given the function k(x)=x^2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: k(x+c) k(x)+c k(cx) ck(x)

OpenStudy (therealmeeeee):

if the original function was k(x) = x^2 then k(x)=x2k(x)+c⟹x2+cvertical shift upwards by "c"k(cx)⟹(c⋅x)2⟹c2⋅x2vertical shrink by a factor of c2c⋅k(x)⟹c⋅x2vertical shrink by a factor of "c"

OpenStudy (anonymous):

uhhh....i dont get it

OpenStudy (ranga):

k(x) = x^2 k(x+c) is replacing x by (x+c). This is equivalent to shifting k(x) to the LEFT by c units. Replacing x by x -c will be shifting k(x) to the right by c units.

OpenStudy (ranga):

In the discussion above and below it is assumed c is positive. k(x)+c shifts k(x) UP by c units. k(x) - c shifts k(x) down by c units. k(cx) stretches or compresses k(x) horizontally. if c is a fraction, that is 0 < c < 1, then it is a horizontal stretching. k(cx) will look wider compared to k(x) when c is a fraction. if c > 1, it is a horizontal compression. k(cx) will look narrower compared to k(x) when c > 1.

OpenStudy (ranga):

ck(x) does vertical stretching or compression. if c is a fraction, that is 0 <c < 1 the graph is compressed vertically If c > 1, the graph is stretched vertically.

OpenStudy (anonymous):

Thank you @ranga